# Spectral Broadening of Landau Levels by a Penetrable Circular Wall

### v1

@@ 1 Introduction
Among the most natural perturbations are singular interactions supported on lower-dimensional sets. $\delta$–type interactions provide models of thin barriers or penetrable–walls and admit a rigorous operator-theoretic treatment[1, 2, 13]. Even without a magnetic field, such interactions can produce curvature-induced bound states [11, 12]. In the magnetic case, $\delta$–interactions supported on general curves were shown to break the degeneracy of Landau levels and to produce discrete eigenvalues accumulating near them [3, 4, 5]. These results, however, are mainly qualitative: they establish the existence of clusters but do not give a detailed description of their internal structure.
@@
Novelty compared with previous work. The general theory developed in [3, 4, 5] shows the existence of eigenvalue clusters near Landau levels for $\delta$–interactions supported on curves of arbitrary shape, but the results are essentially qualitative. By focusing on the circular geometry, we obtain explicit and quantitative information on the fine structure of the broadened spectrum:
@@ 2.4 Penetrable–wall Hamiltonians and gauge covariance
We note that the construction is invariant under gauge transformations. If $A^{\prime}=A+\nabla\chi$ with $\chi$ smooth and real–valued, then the unitary operator $U_{\chi}u=e^{i\chi}u$ satisfies $U_{\chi}H_{B,\alpha}U_{\chi}^{-1}=H_{B,\alpha}$, while the boundary maps transform covariantly (see, e.g., [17, Sec. 15] for gauge covariance in the Schrödinger setting). Indeed, for $U_{\chi}u=e^{i\chi}u$ one has
@@ 3.1 Persistence of the essential spectrum
is a compact embedding. This is a standard Sobolev result; see, for instance, [23, Ch. 4] or classical references on Sobolev spaces. Therefore the composition
@@ 3.2 Behavior of negative eigenvalues for small magnetic fields
The negative eigenvalues of $H_{0,\alpha}$ remain well below the lowest accumulation point $\Lambda_{0}=B$ for small $B$, and hence persist as isolated eigenvalues of $H_{B,\alpha}$. More precisely, Kato’s perturbation theory for analytic families of type (B) (see Kato [16, Chap. VII, Section 3.5]) ensures that every simple eigenvalue $E_{j}(0)<0$ of $H_{0,\alpha}$ generates an analytic eigenvalue branch $E_{j}(B)$ of $H_{B,\alpha}$ with $E_{j}(B)\to E_{j}(0)$ as $B\to 0$. Indeed, with form domain $\mathrm{dom}\,q_{B,\alpha}=H^{1}(\mathbb{R}^{2})$ the closed, lower–bounded quadratic form
@@
depends analytically on $B$. Hence $\{H_{B,\alpha}\}_{B\geq 0}$ is an analytic family of type (B) in the sense of Kato [16, Chap. VII, Section 3.5].
Moreover, by time-reversal symmetry the spectrum is invariant under $B\mapsto-B$, so the eigenvalue branches $E_{j}(B)$ are even functions of $B$ and the linear term in the expansion vanishes. The diamagnetic inequality [21, Eq. (1.1), Chap. 1] implies $E_{j}(B)\geq E_{j}(0)$, hence the quadratic coefficient is nonnegative. Therefore
@@ 3.3 Spectral broadening and direction of accumulation
Although the Landau levels remain as essential spectral accumulation points, the penetrable–wall produces new families of discrete eigenvalues. These eigenvalues are not isolated once and for all, but rather they appear as sequences that converge to the Landau levels. Such broadening phenomena near Landau accumulation points are in line with the general perturbative picture (cf. [20, 18, 9]) and, for $\delta$–type interactions supported on curves, have been analyzed rigorously in [3, 4, 5].
@@
Assume that the wall is the circle $S_{a}$ and $\alpha\in\mathbb{R}\setminus\{0\}$. Then for every $n\in\mathbb{N}_{0}$ there exists an infinite sequence of eigenvalues of $H_{B,\alpha}$, obtained as solutions of the eigenvalue equations of Lemma 4, accumulating at $\Lambda_{n}$. The accumulation is one–sided: if $\alpha<0$ the eigenvalues approach $\Lambda_{n}$ from above, while if $\alpha>0$ they approach from below.
@@
Thus infinitely many $m$ contribute solutions converging to $\Lambda_{n}$, and the sign of the shift shows that the accumulation is from above for $\alpha<0$ and from below for $\alpha>0$. In particular, the eigenvalues originate from the upper gap $(\Lambda_{n},\Lambda_{n+1})$ when $\alpha<0$, and from the lower gap $(\Lambda_{n-1},\Lambda_{n})$ when $\alpha>0$. $\Box$
@@ 4 Explicit form of the Weyl function
To identify $\mu_{m,B}(z)$ we use the spectral resolution of $R_{B}(z)$. Let $\{\psi_{n,k}\}_{n\in\mathbb{N}_{0},\ k\in\mathbb{Z}}$ be an orthonormal basis of Landau eigenfunctions in the symmetric gauge (cf. [17, Sec. 15]), so that
@@ 5 Asymptotic behavior of boundary coefficients
and inserting this in the definition of $|C_{n,m}|^{2}$ yields the claimed asymptotics. The one–sided bound follows from the two–sided Stirling estimates by absorbing constants into $K_{1}$ (see, e.g., [10, Chap. 1] for a convenient reference). $\Box$
@@
together with the elementary binomial estimate $\binom{n+k}{n-j}=\binom{n+k}{n}\frac{n!}{(n-j)!}\,k^{-j}\big{(}1+O(k^{-1})\big{)}$ for fixed $n$, we obtain the stated asymptotics; see, e.g., [10, Chap. 10]. The bound follows by taking absolute values for large $k$ and absorbing constants. $\Box$
@@
with $K_{n}(a,B)=n!Be^{-\tfrac{B}{2}a^{2}}/2$. Finally, using $\binom{n+|m|}{n}\sim(n+|m|)^{n}/n!$ for fixed $n$ and Stirling’s formula for $(n+|m|)!$, we obtain the stated faster than any exponential decay (see also [10, Chap. 10]). $\Box$

### v2

@@ 1 Introduction
Among the most natural perturbations are singular interactions supported on lower-dimensional sets. $\delta$–type interactions provide models of thin barriers or penetrable–walls and admit a rigorous operator-theoretic treatment[1, 2, 11]. Even without a magnetic field, such interactions can produce curvature-induced bound states [9, 10]. In the magnetic case, $\delta$–interactions supported on general curves were shown to break the degeneracy of Landau levels and to produce discrete eigenvalues accumulating near them [3]. These results, however, are mainly qualitative: they establish the existence of clusters but do not give a detailed description of their internal structure.
@@
Novelty compared with previous work. The general theory developed in [3] shows the existence of eigenvalue clusters near Landau levels for $\delta$–interactions supported on curves of arbitrary shape, but the results are essentially qualitative. By focusing on the circular geometry, we obtain explicit and quantitative information on the fine structure of the broadened spectrum:
@@ 2.4 Penetrable–wall Hamiltonians and gauge covariance
We note that the construction is invariant under gauge transformations. If $A^{\prime}=A+\nabla\chi$ with $\chi$ smooth and real–valued, then the unitary operator $U_{\chi}u=e^{i\chi}u$ satisfies $U_{\chi}H_{B,\alpha}U_{\chi}^{-1}=H_{B,\alpha}$, while the boundary maps transform covariantly (see, e.g., [15, Sec. 15] for gauge covariance in the Schrödinger setting). Indeed, for $U_{\chi}u=e^{i\chi}u$ one has
@@ 3.1 Persistence of the essential spectrum
is a compact embedding. This is a standard Sobolev result; see, for instance, [21, Ch. 4] or classical references on Sobolev spaces. Therefore the composition
@@ 3.2 Behavior of negative eigenvalues for small magnetic fields
The negative eigenvalues of $H_{0,\alpha}$ remain well below the lowest accumulation point $\Lambda_{0}=B$ for small $B$, and hence persist as isolated eigenvalues of $H_{B,\alpha}$. More precisely, Kato’s perturbation theory for analytic families of type (B) (see Kato [14, Chap. VII, Section 3.5]) ensures that every simple eigenvalue $E_{j}(0)<0$ of $H_{0,\alpha}$ generates an analytic eigenvalue branch $E_{j}(B)$ of $H_{B,\alpha}$ with $E_{j}(B)\to E_{j}(0)$ as $B\to 0$. Indeed, with form domain $\mathrm{dom}\,q_{B,\alpha}=H^{1}(\mathbb{R}^{2})$ the closed, lower–bounded quadratic form
@@
depends analytically on $B$. Hence $\{H_{B,\alpha}\}_{B\geq 0}$ is an analytic family of type (B) in the sense of Kato [14, Chap. VII, Section 3.5].
Moreover, by time-reversal symmetry the spectrum is invariant under $B\mapsto-B$, so the eigenvalue branches $E_{j}(B)$ are even functions of $B$ and the linear term in the expansion vanishes. The diamagnetic inequality [19, Eq. (1.1), Chap. 1] implies $E_{j}(B)\geq E_{j}(0)$, hence the quadratic coefficient is nonnegative. Therefore
@@ 3.3 Spectral broadening and direction of accumulation
Although the Landau levels remain as essential spectral accumulation points, the penetrable–wall produces new families of discrete eigenvalues. These eigenvalues are not isolated once and for all, but rather they appear as sequences that converge to the Landau levels. Such broadening phenomena near Landau accumulation points are in line with the general perturbative picture (cf. [18, 16, 7]) and, for $\delta$–type interactions supported on curves, have been analyzed rigorously in [3].
@@
Assume that the wall is the circle $S_{a}$ and $\alpha\in\mathbb{R}\setminus\{0\}$. Then for every $n\in\mathbb{N}_{0}$ there exists an infinite sequence of eigenvalues of $H_{B,\alpha}$, obtained as solutions of the eigenvalue equations of Lemma 4, accumulating at $\Lambda_{n}$, except in the case $n=0$ with $\alpha>0$, where no such accumulation occurs. The accumulation is one–sided: if $\alpha<0$ the eigenvalues approach $\Lambda_{n}$ from above, while if $\alpha>0$ they approach from below (for $n=0$ this lower gap is absent).
@@
Thus infinitely many $m$ contribute solutions converging to $\Lambda_{n}$, and the sign of the shift shows that the accumulation is from above for $\alpha<0$ and from below for $\alpha>0$. In particular, for $n\geq 1$ the eigenvalues originate from the upper gap $(\Lambda_{n},\Lambda_{n+1})$ when $\alpha<0$, and from the lower gap $(\Lambda_{n-1},\Lambda_{n})$ when $\alpha>0$. For $n=0$, only the case $\alpha<0$ yields eigenvalues converging to $\Lambda_{0}$ from above. $\Box$
@@ 4 Explicit form of the Weyl function
To identify $\mu_{m,B}(z)$ we use the spectral resolution of $R_{B}(z)$. Let $\{\psi_{n,k}\}_{n\in\mathbb{N}_{0},\ k\in\mathbb{Z}}$ be an orthonormal basis of Landau eigenfunctions in the symmetric gauge (cf. [15, Sec. 15]), so that
@@ 5 Asymptotic behavior of boundary coefficients
and inserting this in the definition of $|C_{n,m}|^{2}$ yields the claimed asymptotics. The one–sided bound follows from the two–sided Stirling estimates by absorbing constants into $K_{1}$ (see, e.g., [8, Chap. 1] for a convenient reference). $\Box$
@@
together with the elementary binomial estimate $\binom{n+k}{n-j}=\binom{n+k}{n}\frac{n!}{(n-j)!}\,k^{-j}\big{(}1+O(k^{-1})\big{)}$ for fixed $n$, we obtain the stated asymptotics; see, e.g., [8, Chap. 10]. The bound follows by taking absolute values for large $k$ and absorbing constants. $\Box$
@@
with $K_{n}(a,B)=n!Be^{-\tfrac{B}{2}a^{2}}/2$. Finally, using $\binom{n+|m|}{n}\sim(n+|m|)^{n}/n!$ for fixed $n$ and Stirling’s formula for $(n+|m|)!$, we obtain the stated faster than any exponential decay (see also [8, Chap. 10]). $\Box$